OpenStudy (anonymous):

solve the system: 2x-4y=-10, 3x +4y=-5

OpenStudy (anonymous):

2x-4y=-10, 3x+4y=-5 2(3x+4y)=2(-5) We multiply the second equation by two getting a similar equation. 2x-4y=-10 6x+8y=-10 We have the same system, now we only have to sum the corresponding parts 6x+8y=-10 take out the value of X?